A block with mass m =7.2 kg is hung from a vertica
Answer #1

mass m = 7.2 kg

Streachness X = 0.25 m

Velocity at equilibrium position Vo = 4.6 m/s

(1). At equlibrium , mg = kX

from this spring constant k = mg / X

k = (7.2)(9.8)/(0.25)

= 282.24 N/m

(2).Oscillation frequency f = (1/2\pi) \sqrt{}[k/m]

= (1/6.28) \sqrt{}[282.24/7.2]

= 0.9964 Hz

(3).Velocity of the block v(t) = Vo cos (wt) where w = angular frequency = 2(pi) f = 6.26 rad/s

= 4.6 cos (6.26 t)

velocity at t= 0.39 s is v(0.39) = 4.6 cos(6.26x0.39 rad)

= 4.6 (-0.7649)

= -3.518 m/s

(d). Maximum acceleration = Aw 2

= (Aw)w

= Vo w

= 4.6(6.26)

= 28.8 m/s 2